Your company is considering a machine that will cost $ 3,813 at Time 0 and which can be sold after 3 years for $ 321 . To operate the machine, $ 486 must be invested at Time 0 in inventories; these funds will be recovered when the machine is retired at the end of Year 3. The machine will produce sales revenues of $ 826 /year for 3 years; variable operating costs (excluding depreciation) will be 55 percent of sales. Operating cash inflows will begin 1 year from today (at Time 1). The machine is in the 3-year MACRS class. The MACRS class has depreciation of 33% in year 1, 45% in year 2, 15% in year 3, and 7% in year 4. The company has a 33 percent tax rate, enough taxable income from other assets to enable it to get a tax refund from this project if the project's income is negative, and a 10 percent cost of capital. Inflation is zero. What are the terminal cash flows associated with ending this project?
Answer #1

Terminal cash flow = After tax salvage value of machine + Recovery of working capital investment

Book value after 3 years = 3813*(100-33-45-15)% = $266.91

Selling price = $321

Gain on Sale = $54.09

Tax on gain = 54.09*33% = $17.8497

After tax salvage value = 321-17.8497 = $303.1503

Hence, terminal cash flows = 303.1503 + 486 = $789.1503